Montserrat Volcano Observatory

Daily Report
Report for the period 6 pm 12 February
to 6 pm 13 February 1998

Activity at the volcano was about the same as yesterday with dust levels generally low and rockfall events dominant.

Between 16:00 yesterday and 16:00 today there were 7 hybrid earthquakes, 9 long period earthquakes, 2 volcano-tectonic earthquakes and 91 distinct rockfall signals. There were three periods of heightened activity during which most of the rockfalls occurred. These were experienced between 6 pm and 8 pm last night; between 1 am and 4 am and between 11 am to 1 pm today.

Ash and steam plumes generated by ash venting and rockfall activity were observed at the volcano throughout the day. Rockfalls continue to occur on the new dome which now fills the December 26th collapse scar.

The Harris and Hermitage GPS stations were occupied during the day with a long occupation of Blakes and Harris currently underway. EDM measurements were made to the reflector at Lees Yard today. The results of these measurements are currently being processed.

A new electronic tilt station was installed at Hermitage today and the signal relayed directly to the MVO at Mongo Hill. The data is already being collected but it may be a few more days before any clear trends can be determined.

Dust levels at all except one of the monitoring sites remained low today. One site in Woodlands was raised.

People should remain alert and listen to Radio Montserrat for further information.

Montserrat Volcano Observatory